OpenStudy (anonymous):

Jana has a rectangular sheet of cardboard that measures 10 inches by 18 inches. What is themaximum number of 3-inch by 3-inch squares that she can cut from this sheet of cardboard?

jigglypuff314 (jigglypuff314):

first find the area of the rectangle (A = L * W) then divide it by area of square pieces. so 10 * 18 = ___ 3 * 3 = 9 then (10*18)/9 = how many squares can be cut from this rectangle